OpenStudy (anonymous):

Choose the correct slope of the line shown on the graph below. graph of a line passing through points negative 5 comma negative 2 and 5 comma 6 −four fifths −five fourths five fourths four fifths

OpenStudy (anonymous):

@e.mccormick ? help

OpenStudy (e.mccormick):

Wait, the x values are the same and those are the only answer they give?

OpenStudy (e.mccormick):

OH! Missed a negative. DOH!

OpenStudy (e.mccormick):

OK, did you put it into the formula I talked about on the last one?

OpenStudy (anonymous):

i tried im still confused cause its a timed quiz so im trynna answer multiple questions at once.

OpenStudy (anonymous):

ohk do u know slope formula @InstagramDoll

OpenStudy (e.mccormick):

OK, so it was \[\frac{y_1-y_2}{x_1-x_2}\]What would that be in this case?

OpenStudy (anonymous):

my bad my computer posted that acting up lol

OpenStudy (anonymous):

(-5,-2); (5,6) here (x,y) slope = \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (anonymous):

now simply replace the values in x and y's

OpenStudy (anonymous):

got it or still confused?

OpenStudy (e.mccormick):

It is also good to note that it does not matter which point you choose as \((x_1,y_1)\) or \((x_2,y_2)\). Either way gets the same answer! So: \((x_1,y_1)=(-5,-2)\) and \((x_2,y_2)=(5,6)\) OR \((x_1,y_1)=(5,6)\) and \((x_2,y_2)=(-5,-2)\) You choose and get the same answer either way.

OpenStudy (anonymous):

so its -4 / 5

OpenStudy (anonymous):

@e.mccormick @Best_Mathematician

OpenStudy (e.mccormick):

Not negative. It would be -8/-10 or 8/10, so the - would cancel.

OpenStudy (anonymous):

-4/5 bc subtract 6-2 then 5-5 = o then it cancels bt the 5 comes back sum how giving u tat answer

OpenStudy (anonymous):

thats how i did it..

OpenStudy (e.mccormick):

\[\frac{6-(-2)}{(5-(-5))}=\frac{6+2}{5+5}\]

OpenStudy (anonymous):

so basically the answer will be 4/5 since the - cancels out?

OpenStudy (e.mccormick):

Yah. And it would cancel the other way too.

OpenStudy (anonymous):

thanks

OpenStudy (e.mccormick):

\[\frac{(-2)-6}{(-5)-5}=\frac{-8}{-10}\]So they both are - and that cances, so same answer. I hope that is clearer now.

OpenStudy (anonymous):

Wait so I was right when I said -4/5 or its 4/5